# Two-Photon Absorption in Silicon Using Real Density Matrix Approach

### v1

@@ II.2 Discrete states
In section II.4, we use the expressions for the nonlinear susceptibility above to determine the nonlinear absorption coefficients in silicon.
II.3 Continuum states
If $(\hbar\omega-\hbar\omega_{ph})>E_{g}$ for the case of phonon emission, or $(\hbar\omega+\hbar\omega_{ph})>E_{g}$ for the case of phonon absorption, then there is no generation of bound exciton states, and continuum states constitute the final states instead. In this case, $Y^{(1)}$ is calculated in the first iteration step by setting $V_{eh}=0$ in the electron-hole Hamiltonian [24], giving rise to equations of the form
$\left(E_{g}\pm\hbar\omega\pm\hbar\omega_{ph}-i{\mit\Gamma}-\frac{\hbar^{2}}{2\mu}\hbox{\boldmath$\nabla$}^{2}\right)Y_{\pm}=\textbf{M}(\textbf{r})\textbf{E}.$ (39)
Equation (39) can be solved by means of the appropriate Green function
$Y_{\pm}^{(1)}=\int d^{3}\,r^{\prime}g_{\pm}(r,r^{\prime})\textbf{M}(r^{\prime},\theta,\phi)\textbf{E},$ (40)
@@ II.3 Continuum states
$g_{\pm}(r,r^{\prime})=\frac{2\mu}{\hbar^{2}}\frac{\sinh\kappa_{\pm}r^{<}}{4\pi\kappa_{\pm}r^{<}r^{>}}e^{-\kappa_{\pm}r^{>}},$ (41)
@@
$\kappa^{2}_{\pm}=\frac{2\mu}{\hbar^{2}}\left(E_{g}\pm\hbar\omega\pm\hbar\omega_{ph}-i{\mit\Gamma}\right).$ (42)
Assuming a linear polarization and the wave vector E having a component $E_{0}$ in a direction $\alpha$, simultaneously with the dipole density M having a component $M_{0}$ in the same direction and, for simplicity, using the dipole density of the form
$M(\textbf{r})=\frac{M_{0}\delta(r-r_{0})}{4\pi r_{0}^{2}}e^{-r/r_{0}}$
@@
$Y^{(1)}_{\pm}=M_{0}E_{0}\,g_{\pm}(r,r_{0}).$ (43)
We may again define amplitudes of the form $Y^{(1)}_{ca\pm},Y^{(1)}_{cb\pm}$, where the subscript “c” now indicates the involvement of continuum states. If, in the case of phonon emission, we encounter $\kappa^{2}<0$, we introduce
$\kappa=-i\tilde{\kappa},$ (44)
@@
$\tilde{\kappa}^{2}_{-}=\hbar\omega-\hbar\omega_{ph}-E_{g}+i\Gamma,$
@@
$g_{-}(r,r^{\prime})=\frac{2\mu}{\hbar^{2}}\frac{\sin\tilde{\kappa}_{\pm}r^{<}}{4\pi\tilde{\kappa}_{\pm}r^{<}r^{>}}e^{i\tilde{\kappa}_{\pm}r^{>}}.$ (45)
@@
$g_{a-,em}(r,r_{0})=\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\sin(\tilde{\kappa}_{a-,em}r^{<})}{4\pi\tilde{\kappa}_{a-,em}rr_{0}}e^{i\tilde{\kappa}_{a-,em}r^{>}},$
$\tilde{\kappa}_{a-,em}^{2}=\frac{2\mu_{\parallel H}}{\hbar^{2}}\left[\hbar\omega_{a}-(E_{g}+\hbar\omega_{ph})\right]+i\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\hbar}{T_{2}}.$
Similar expressions can be obtained for the amplitudes at input frequencies $\omega_{b}$. The linear amplitudes thus obtained form the source for calculating the $C$ and $D$ matrices, followed by third step to determine $Y^{(3)}_{ca,b\pm}$, similar to the procedure described in section II.2. Once the $Y^{(3)}_{ca,b\pm}$ amplitudes are found, for both the phonon emission and absorption process, we may write the nonlinear cross susceptibility for the continuum states can as
$\chi^{(3)}_{\hbox{\tiny c,cross}}=\chi^{(3)}_{\hbox{\tiny c,cross,em}}(\omega_{a},\omega_{b})+\chi^{(3)}_{\hbox{\tiny c,cross,abs}}(\omega_{a},\omega_{b}),$ (47)
@@
$\chi^{(3)}_{\hbox{\tiny c,cross,em}}(\omega_{a},\omega_{b})=$
$=-(n_{ph}+1)\frac{2}{\epsilon_{0}}2\frac{T_{1}}{\hbar}M_{0}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}$
×$\times\frac{1}{4\pi r_{0}}\left(\frac{\sin(\tilde{\kappa}_{a-,em}r_{0})}{\tilde{\kappa}_{a-,em}r_{0}}\right)^{2}$
×$\times[\tilde{\kappa}_{b-,em}r_{0}+\tilde{\kappa}_{b-,abs}r_{0}](A_{e}+A_{h}),$ (48)
$\chi^{(3)}_{\hbox{\tiny c,cross,abs}}(\omega_{a},\omega_{b})=$
$=-n_{ph}\frac{2}{\epsilon_{0}}2\frac{T_{1}}{\hbar}M_{0}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}$
×$\times\frac{1}{4\pi r_{0}}\left(\frac{\sin(\tilde{\kappa}_{a-,abs}r_{0})}{\tilde{\kappa}_{a-,abs}r_{0}}\right)^{2}$
×$\times[\tilde{\kappa}_{b-,em}r_{0}+\tilde{\kappa}_{b-,abs}r_{0}](A_{e}+A_{h}),$ (49)
@@
$\tilde{\kappa}_{j-,em}r_{0}=\left(x_{j}-\frac{E_{g}+\hbar\omega_{ph}}{E_{g}}\right)^{1/2},$
$\tilde{\kappa}_{j-,abs}r_{0}=\left(x_{j}-\frac{E_{g}-\hbar\omega_{ph}}{E_{g}}\right)^{1/2},$
Here $A_{e},A_{h}$ correspond to the expressions (II.2), related to the irreversible processes
$A_{e}$ $=$ $A_{e}^{\prime}+iA_{e}^{\prime\prime}=\int d^{3}r^{\prime}g_{a-}(r_{0},r^{\prime})f_{0e}(r^{\prime})$
[…]
with an analogous expression for $A_{h}$.
[…]
Parameter
Value
Unit
Reference
$E_{g}$(4.2K)
1170
meV
[30]
$E_{g}$(300K)
1124
meV
[30]
$R_{H}^{*}$(4.2K)
13.61
meV
$R_{H}^{*}$(300  K)
15.65
meV
$R_{L}^{*}$(0  K)
8.64
meV
$R_{L}^{*}$(300  K)
8.35
meV
$\Delta_{LTH}$
$0.1$
meV
[31]
$m_{e,\hbox{\tiny\emph{long}}}(4.2\hbox{K})$
0.9163
$m_{0}$
[32]
$m_{e,\hbox{\tiny\emph{long}}}(300\hbox{K})$
1.09 $m_{0}$
[30]
$m_{e,\hbox{\tiny\emph{tr}}}$
0.1905
$m_{0}$
[32]
$m_{e,\hbox{\tiny\emph{dos}}}(4.2\hbox{K})$
1.04
$m_{0}$
$m_{e,\hbox{\tiny\emph{dos}}}(300\hbox{K})$
1.12
$m_{0}$
$m_{hH}$
0.49
$m_{0}$
[33]
$m_{hL}$
0.16
$m_{0}$
[33]
$m_{hH}(4.2\hbox{K})$
0.59
$m_{0}$
[30]
$m_{hH}(300\hbox{K})$
1.15
$m_{0}$
[30]
$\mu_{H,\hbox{\tiny\emph{long}}}(4.2\hbox{K})$
0.319
$m_{0}$
$\mu_{H,\hbox{\tiny\emph{long}}}(300\hbox{K})$
0.56
$m_{0}$
$\mu_{H,\hbox{\tiny\emph{tr}}}(4.2\hbox{K})$
0.137
$m_{0}$
$\mu_{H,\hbox{\tiny\emph{tr}}}(300\hbox{K})$
0.163
$m_{0}$
$\mu_{L,\hbox{\tiny\emph{long}}}(4.2\hbox{K})$
0.136
$m_{0}$
$\mu_{L,\hbox{\tiny\emph{long}}}(300\hbox{K})$
0.14
$m_{0}$
$\mu_{L,\hbox{\tiny\emph{tr}}}(4.2\hbox{K})$
0.087
$m_{0}$
$\mu_{L,\hbox{\tiny\emph{tr}}}(300\hbox{K})$
0.087
$m_{0}$
$a^{*}_{H}$(4.2 K)
4.52
 nm
[…]
3.86
 nm
$a^{*}_{L}$(4.2 K)
7.11
 nm
$a^{*}_{L}$(300 K)
7.235
 nm
$r_{0H}(4.2\hbox{K})$
0.487
nm
$r_{0H}(300\hbox{K})$
0.455
nm
$r_{0L}(4.2\hbox{K})$
0.61
nm
$r_{0L}(300\hbox{K})$
0.61
nm
$\epsilon_{b}$(4.2K)
11.7
$\epsilon_{b}$(300K)
11.9
$T_{2}$
0.1
ns
[26]
$T_{1}$
8
ns
fitting
$\hbar\omega_{ph}$
40
meV
fitting, [28]
II.4 Nonlinear absorption coefficients
@@
$\frac{\partial E_{0a}}{\partial z}=i\frac{\omega_{a}}{n_{a}\,c}\chi^{(3)}_{\hbox{\tiny self}}(\omega_{a},\omega_{a})|E_{0a}|^{2}\,E_{0a}$ (50)
$+i\frac{\omega_{a}}{n_{a}\,c}\chi^{(3)}_{\hbox{\tiny cross}}(\omega_{a},\omega_{b})|E_{0b}|^{2}\,E_{0a}\;,$
$\frac{\partial E_{0b}}{\partial z}=i\frac{\omega_{b}}{n_{b}\,c}\chi^{(3)}_{\hbox{\tiny self}}(\omega_{b},\omega_{b})|E_{0b}|^{2}\,E_{0b}$ (51)
$+i\frac{\omega_{b}}{n_{b}\,c}\chi^{(3)}_{\hbox{\tiny cross}}(\omega_{b},\omega_{a})|E_{0a}|^{2}\,E_{0b}\;.$
@@
$\alpha_{2}(\omega_{a},\omega_{a})=\frac{\omega_{a}}{2\epsilon_{0}\,n_{a}^{2}c^{2}}\,\chi^{(3),I}_{\hbox{\tiny self}}(\omega_{a},\omega_{a}),$
$\alpha_{2}(\omega_{a},\omega_{b})=\frac{\omega_{a}}{2\epsilon_{0}\,n_{a}\,n_{b}\,c^{2}}\,\,\chi^{(3),I}_{\hbox{\tiny cross}}(\omega_{a},\omega_{b}),$
$\alpha_{2}(\omega_{b},\omega_{a})=\frac{\omega_{b}}{2\epsilon_{0}\,n_{a}\,n_{b}\,c^{2}}\,\chi^{(3),I}_{\hbox{\tiny cross}}(\omega_{b},\omega_{a}),$
$\alpha_{2}(\omega_{b},\omega_{b})=\frac{\omega_{b}}{2\epsilon_{0}\,n_{b}^{2}c^{2}}\,\,\chi^{(3),I}_{\hbox{\tiny self}}(\omega_{b},\omega_{b}),$
where $\chi^{(3),I}$ indicates the imaginary part of $\chi^{(3)}$. The refractive indices $n_{a}=n(\omega_{a}),n_{b}=n(\omega_{b})$ are defined by the real parts of the linear susceptibility $\chi^{(1)}$, defined as
@@
with analogous formulas for $\omega_{b}$ and the absorptive term $\hbar\Omega_{n,abs}$. Here, $f_{1}$ is the oscillator strength (here for the heavy hole exciton), written as
@@
The susceptibilities and, in consequence, the nonlinear absorption coefficients, are composed of 4 components: two corresponding to the contributions of discrete and continuous states, and each of them containing terms related to phonon emission and absorption
$\alpha_{2}(\omega_{a},\omega_{b})=\alpha_{2d}(\omega_{a},\omega_{b})+\alpha_{2c}(\omega_{a},\omega_{b}),$ (57)
@@
$\alpha_{2d}(\omega_{a},\omega_{b})$
$=-2{x_{a}\gamma_{2}\alpha^{\prime}}\left(\frac{T_{1}}{T_{2}}\right)\biggl{\{}\frac{\epsilon_{b}}{n_{a}n_{b}}\frac{n_{ph}+1}{(\hbar\Omega_{1,em}/E_{g}-x_{b})^{2}+\gamma_{2}^{2}}$
×$\times\frac{\hbar\Omega_{1,em}/E_{g}}{[(\hbar\Omega_{1,em}/E_{g})^{2}-x_{a}^{2}]^{2}+(2\gamma_{2}x_{a})^{2}}$
$+\frac{\epsilon_{b}}{n_{a}n_{b}}\frac{n_{ph}}{(\hbar\Omega_{1,abs}/E_{g}-x_{a})^{2}+\gamma_{2}^{2}}$
×$\times\frac{\hbar\Omega_{1,abs}/E_{g}}{[(\hbar\Omega_{1,abs}/E_{g})^{2}-x_{b}^{2}]^{2}+(2\gamma_{2}x_{b})^{2}}\biggr{\}},$ (58)
@@
$\alpha^{\prime}(\omega_{a},\omega_{b})$
$=2\frac{4E_{g}}{\epsilon_{0}\hbar n_{a}\,n_{b}2c^{2}E_{g}^{3}}\frac{2M_{0}^{2}}{\epsilon_{0}}\left[\varphi_{1H}(0)A_{1}+\varphi_{1H}(0)B_{1}\right]c_{1H}^{2}$
@@
×$\times\left[\varphi_{1H}(0)A_{1}+\varphi_{1H}(0)B_{1}\right]\left(\frac{\eta_{H}}{a^{*}_{H}}\right)^{3}\frac{1}{(1+\eta_{H}r_{0H}/a^{*}_{H})^{2}}$
$=2\frac{\pi\epsilon_{b}^{2}\Delta_{LTH}^{2}(\eta_{H}a^{*}_{H})^{3}}{E_{g}^{2}\hbar c^{2}\,n_{a}\,n_{b}}\frac{\left[\varphi_{1H}(0)A_{1}+\varphi_{1H}(0)B_{1}\right]}{(1+\eta_{H}r_{0H}/a^{*}_{H})^{2}}.$
@@
$\alpha_{2c}(\omega_{a},\omega_{b})=\alpha_{2c,em}(\omega_{a},\omega_{b})+\alpha_{2c,abs}(\omega_{a},\omega_{b}),$ (59)

### v2

@@ II.4 Discrete states - Nonlinear susceptibility
In section II.6, we use the expressions for the nonlinear susceptibility above to determine the nonlinear absorption coefficients in silicon.
II.5 Continuum states
In this section, we consider the case relevant to heavy-hole exciton transitions. If $(\hbar\omega-\hbar\omega_{ph})>E_{g}$ for the case of phonon emission, or $(\hbar\omega+\hbar\omega_{ph})>E_{g}$ for the case of phonon absorption, then there is no generation of bound exciton states, and continuum states constitute the final states instead. In this case, $Y^{(1)}_{H}$ is calculated in the first iteration step by setting $V_{12}=0$ in the electron-hole Hamiltonian [24], giving rise to equations of the form
$\left(E_{g}\pm\hbar\omega\pm\hbar\omega_{ph}-i{\frac{\hbar}{T_{2}}}-\frac{\hbar^{2}}{2\mu_{\parallel H}}\hbox{\boldmath$\nabla$}^{2}\right)Y_{H\pm}$
$=\textbf{M}_{H}(\textbf{r})\textbf{E}.$ (81)
Equation (II.5) can be solved by means of the appropriate Green function
$Y_{H\pm}^{(1)}=\int d^{3}\,r^{\prime}g_{H\pm}(r,r^{\prime})\textbf{M}_{H}(r^{\prime},\theta,\phi)\textbf{E},$ (82)
@@ II.5 Continuum states
$g_{H\pm}(r,r^{\prime})=\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\sinh\kappa_{H\pm}r^{<}}{4\pi\kappa_{H\pm}r^{<}r^{>}}e^{-\kappa_{H\pm}r^{>}},$ (83)
@@
$\kappa^{2}_{H\pm}=\frac{2\mu_{\parallel H}}{\hbar^{2}}\left(E_{g}\pm\hbar\omega\pm\hbar\omega_{ph}-i\frac{\hbar}{T_{2}}\right).$ (84)
Assuming a linear polarization and the wave vector E having a component $E_{0}$ in a direction $\alpha$, simultaneously with the dipole density $\textbf{M}_{H}$ having a component $M_{0H}$ in the same direction and, for simplicity, using the dipole density of the form
$M_{H}(\textbf{r})=\frac{M_{0H}\delta(r-r_{0H})}{4\pi r_{0H}^{2}},$ (85)
@@
$Y^{(1)}_{H\pm}=M_{0H}E_{0}\,g_{H\pm}(r,r_{0H}).$ (86)
We may again define amplitudes of the form $Y^{(1)}_{caH\pm},Y^{(1)}_{cbH\pm}$, where the subscript “c” now indicates the involvement of continuum states. If, in the case of phonon emission, we encounter $\kappa_{H}^{2}<0$, and introduce
$\kappa_{H}=-i\tilde{\kappa_{H}},$ (87)
@@
$\tilde{\kappa}^{2}_{H-}=\frac{2\mu_{\parallel H}}{\hbar^{2}}(\hbar\omega-\hbar\omega_{ph}-E_{g}+i\frac{\hbar}{T_{2}}),$
@@
$g_{H-}(r,r^{\prime})=\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\sin\tilde{\kappa}_{H\pm}r^{<}}{4\pi\tilde{\kappa}_{H\pm}r^{<}r^{>}}e^{i\tilde{\kappa}_{H\pm}r^{>}}.$ (88)
@@
$g_{aH-,em}(r,r_{0H})=\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\sin(\tilde{\kappa}_{aH-,em}r^{<})}{4\pi\tilde{\kappa}_{aH-,em}rr_{0}}e^{i\tilde{\kappa}_{aH-,em}r^{>}},$
$\tilde{\kappa}_{aH-,em}^{2}=\frac{2\mu_{\parallel H}}{\hbar^{2}}\left[\hbar\omega_{a}-(E_{g}+\hbar\omega_{ph})\right]+i\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\hbar}{T_{2}}.$
Similar expressions can be obtained for the amplitudes at input frequencies $\omega_{b}$. The linear amplitudes thus obtained form the source for calculating the $C_{H}$ and $D_{H}$ matrices, followed by third step to determine $Y^{(3)}_{ca,bH\pm}$, similar to the procedure described in section II.3. Once the $Y^{(3)}_{ca,bH\pm}$ amplitudes are found, for both the phonon emission and absorption process, we may write the nonlinear cross susceptibility for the continuum states as
$\chi^{(3)}_{\hbox{\tiny cH,cross}}=\chi^{(3)}_{\hbox{\tiny cH,cross,em}}(\omega_{a},\omega_{b})+\chi^{(3)}_{\hbox{\tiny cH,cross,abs}}(\omega_{a},\omega_{b}),$ (90)
@@
$\chi^{(3)}_{\hbox{\tiny cH,cross,em}}(\omega_{a},\omega_{b})=$ (91)
$=-(n_{ph}+1)\frac{2}{\epsilon_{0}}2\frac{T_{1}}{\hbar}M_{0H}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}$
×$\times\frac{1}{4\pi r_{0H}}\left(\frac{\sin(\tilde{\kappa}_{aH-,em}r_{0H})}{\tilde{\kappa}_{aH-,em}r_{0H}}\right)^{2}$
×$\times[\tilde{\kappa}_{bH-,em}r_{0H}+\tilde{\kappa}_{bH-,abs}r_{0H}]({\mathcal{A}}_{eH,em}+{\mathcal{B}}_{hH,em}),$
$\chi^{(3)}_{\hbox{\tiny cH,cross,abs}}(\omega_{a},\omega_{b})=$ (92)
$=-n_{ph}\frac{2}{\epsilon_{0}}2\frac{T_{1}}{\hbar}M_{0H}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}$
×$\times\frac{1}{4\pi r_{0H}}\left(\frac{\sin(\tilde{\kappa}_{aH-,abs}r_{0H})}{\tilde{\kappa}_{aH-,abs}r_{0H}}\right)^{2}$
×$\times[\tilde{\kappa}_{bH-,em}r_{0H}+\tilde{\kappa}_{bH-,abs}r_{0H}]({\mathcal{A}}_{eH,abs}+{\mathcal{B}}_{hH,abs}),$
@@
$\tilde{\kappa}_{jH-,em}r_{0H}=\left(x_{j}-\frac{E_{g}+\hbar\omega_{ph}}{E_{g}}\right)^{1/2},$
$\tilde{\kappa}_{j-,abs}r_{0H}=\left(x_{j}-\frac{E_{g}-\hbar\omega_{ph}}{E_{g}}\right)^{1/2},$
$x_{j}=\frac{\hbar\omega_{j}}{E_{g}}.$ (93)
Here ${\mathcal{A}}_{eH,em},{\mathcal{B}}_{hH,em},{\mathcal{A}}_{eH,abs},{\mathcal{B}}_{hH,abs}$ are the continuum counterparts to expressions (II.4), appropriate for discrete states. Making use of the relations (II.4), we obtain
${\mathcal{A}}_{eH,em}={\mathcal{A}}_{eH,em}^{\prime}+i{\mathcal{A}}_{eH,em}^{\prime\prime}=\frac{2\mu_{\parallel H}}{\hbar^{2}}\frac{\sin(\tilde{\kappa}_{aH-,em}r_{0H})}{\tilde{\kappa}_{aH-,em}r_{0H}}$
×$\times\int\limits_{\rho_{0H}}^{\infty}\rho\,d\rho\int\limits_{0}^{\infty}dz\frac{\exp(-i\tilde{\kappa}_{aH-,em})\sqrt{\rho^{2}+(z^{2}/\gamma_{aH})}}{\sqrt{\rho^{2}+(z^{2}/\gamma_{aH})}}$
×$\times\exp\left(-\frac{\rho^{2}}{2\lambda_{th,e\parallel}^{2}}-\frac{z^{2}}{2\lambda_{th,ez}^{2}}\right),$ (94)
with an analogous expression for ${\mathcal{A}}_{eH,abs},\;{\mathcal{B}}_{hH,em}$, and ${\mathcal{B}}_{hH,abs}$. The value $\rho_{0H}=r_{0h}/a_{H}^{*}$.
Table 1: Band parameter values for Si, masses in free electron mass $m_{0}$, H denotes heavy-, and L light-hole, $R^{*}_{H,L}$ calculated from ($\mu_{\parallel H,L}/\epsilon_{b}^{2})\times 13600\,\hbox{meV}$,  $a^{*}_{H,L}$ calculated from $(1/\mu_{\parallel H,L})\epsilon_{b}\times 0.0529\,\hbox{nm}$, $m_{e,dos}$ (density-of-state effective mass) calculated from $6^{2/3}(m_{e\parallel}m_{e\parallel}m_{ez})^{1/3}$.
Parameter Value (4.2K) Value(300K) Unit Reference
$E_{g}$ 1170 1124 meV [42]
$\Delta_{LTH}$ 0.1 meV
$R_{H}^{*}$ 15 16.56 meV
$R^{*}_{L}$ 7.94 9.6 meV
$\gamma_{1}$ 4.285 2.45 [43],
$\gamma_{2}$ 0.339 0.194 [43],
$\gamma_{3}$ 1.446 0.826 [43],
$m_{ez}$ 0.9163 1.09 $m_{0}$ [44]
$m_{e\parallel}$ 0.1905 0.2 $m_{0}$ [44]
$m_{e,dos}$ 1.06 1.16 $m_{0}$
$m_{hzH}$ 0.28 0.485 $m_{0}$ Eq.(II.6),[45]
$m_{h\parallel H}$ 0.72 1.25 $m_{0}$ Eq. (II.6)
$m_{hzL}$ 0.2 0.35 $m_{0}$
$m_{h\parallel L}$ 0.14 0.24 $m_{0}$
$\mu_{zH}$ 0.214 0.336 $m_{0}$
$\mu_{\parallel H}$ 0.15 0.172 $m_{0}$
$\mu_{zL}$ 0.164 0.26 $m_{0}$
$\mu_{\parallel L}$ 0.08 0.1 $m_{0}$
$a^{*}_{H}$ 4.13 3.66 nm
$a^{*}_{L}$ 7.74 6.3 nm
$r_{0H}$ 0.46 0.44 nm
$r_{0L}$ 0.64 0.58 nm
$\epsilon_{b}$ 11.7 11.9
$T_{2}$ 0.1 ns [46]
$T_{1}$ 8 ns fitting
$\hbar\omega_{ph}$ 40 meV fitting, [38]
Table 2: Anisotropy parameters for Si, excitonic energies calculated from Eq. (II.3)
Parameter Value (4.2 K) Value (300 K) Unit Reference
$\gamma_{aH}$ 0.7 0.51 Eq.(55)
$\eta_{00H}$ 1.058 1.1 Eq.(55)
$|E_{10H}|$ 16.79 20 meV
II.6 Nonlinear absorption coefficients
@@
$\frac{\partial E_{0a}}{\partial z}=i\frac{\omega_{a}}{n_{a}\,c}\chi^{(3)}_{\hbox{\tiny self H}}(\omega_{a},\omega_{a})|E_{0a}|^{2}\,E_{0a}$ (95)
$+i\frac{\omega_{a}}{n_{a}\,c}\chi^{(3)}_{\hbox{\tiny cross H}}(\omega_{a},\omega_{b})|E_{0b}|^{2}\,E_{0a}\;,$
$\frac{\partial E_{0b}}{\partial z}=i\frac{\omega_{b}}{n_{b}\,c}\chi^{(3)}_{\hbox{\tiny self H}}(\omega_{b},\omega_{b})|E_{0b}|^{2}\,E_{0b}$ (96)
$+i\frac{\omega_{b}}{n_{b}\,c}\chi^{(3)}_{\hbox{\tiny cross H}}(\omega_{b},\omega_{a})|E_{0a}|^{2}\,E_{0b}\;.$
@@
$\alpha_{2H}(\omega_{a},\omega_{a})=\frac{\omega_{a}}{2\epsilon_{0}\,n_{a}^{2}c^{2}}\,\chi^{(3),I}_{\hbox{\tiny self H}}(\omega_{a},\omega_{a}),$
$\alpha_{2H}(\omega_{a},\omega_{b})=\frac{\omega_{a}}{2\epsilon_{0}\,n_{a}\,n_{b}\,c^{2}}\,\,\chi^{(3),I}_{\hbox{\tiny cross H}}(\omega_{a},\omega_{b}),$
$\alpha_{2H}(\omega_{b},\omega_{a})=\frac{\omega_{b}}{2\epsilon_{0}\,n_{a}\,n_{b}\,c^{2}}\,\chi^{(3),I}_{\hbox{\tiny cross H}}(\omega_{b},\omega_{a}),$
$\alpha_{2H}(\omega_{b},\omega_{b})=\frac{\omega_{b}}{2\epsilon_{0}\,n_{b}^{2}c^{2}}\,\,\chi^{(3),I}_{\hbox{\tiny self H}}(\omega_{b},\omega_{b}),$
where $\chi_{H}^{(3),I}$ indicates the imaginary part of $\chi_{H}^{(3)}$. The refractive indices $n_{a}=n(\omega_{a}),n_{b}=n(\omega_{b})$ are defined by the real parts of the linear susceptibility $\chi^{(1)}$, defined as
@@
with analogous formulas for $\omega_{b}$ and the absorptive term $\hbar\Omega_{n,abs}$. Here, $\gamma_{2}=\hbar/(E_{g}T_{2})$, $f_{1}$ is the oscillator strength (here for the heavy hole exciton), written as
@@
The susceptibilities and, consequently, the nonlinear absorption coefficients, are composed of four components: two corresponding to the contributions of discrete and continuous states, and each of them containing terms related to phonon emission and absorption
$\alpha_{2H}(\omega_{a},\omega_{b})=\alpha_{2dH}(\omega_{a},\omega_{b})+\alpha_{2cH}(\omega_{a},\omega_{b}),$ (102)
@@
$\alpha_{2dH}(\omega_{a},\omega_{b})$
$=-2{x_{a}\gamma_{2}\alpha^{\prime}}\left(\frac{T_{1}}{T_{2}}\right)\biggl{\{}\frac{\epsilon_{b}}{n_{a}n_{b}}\frac{n_{ph}+1}{(\hbar\Omega_{1H,em}/E_{g}-x_{b})^{2}+\gamma_{2}^{2}}$
×$\times\frac{\hbar\Omega_{1H,em}/E_{g}}{[(\hbar\Omega_{1H,em}/E_{g})^{2}-x_{a}^{2}]^{2}+(2\gamma_{2}x_{a})^{2}}$
$+\frac{\epsilon_{b}}{n_{a}n_{b}}\frac{n_{ph}}{(\hbar\Omega_{1H,abs}/E_{g}-x_{a})^{2}+\gamma_{2}^{2}}$
×$\times\frac{\hbar\Omega_{1H,abs}/E_{g}}{[(\hbar\Omega_{1H,abs}/E_{g})^{2}-x_{b}^{2}]^{2}+(2\gamma_{2}x_{b})^{2}}\biggr{\}},$ (103)
@@
$\alpha_{H}^{\prime}(\omega_{a},\omega_{b})$
$=2\frac{4E_{g}}{\epsilon_{0}\hbar n_{a}\,n_{b}2c^{2}E_{g}^{3}}\frac{2M_{0H}^{2}}{\epsilon_{0}}\left[\varphi_{1H}(0)A_{1H}+\varphi_{1H}(0)B_{1H}\right]c_{1H}^{2}$
@@
×$\times\left[\varphi_{1H}(0)A_{1H}+\varphi_{1H}(0)B_{1H}\right]\left(\frac{\eta_{00H}}{a^{*}_{H}}\right)^{3}\frac{1}{(1+\eta_{00H}r_{0H}/a^{*}_{H})^{2}}$
$=2\frac{\pi\epsilon_{b}^{2}\Delta_{LTH}^{2}(\eta_{00H}a^{*}_{H})^{3}}{E_{g}^{2}\hbar c^{2}\,n_{a}\,n_{b}}\frac{\left[\varphi_{1H}(0)A_{1H}+\varphi_{1H}(0)B_{1H}\right]}{(1+\eta_{00H}r_{0H}/a^{*}_{H})^{2}}.$
@@
$\alpha_{2cH}(\omega_{a},\omega_{b})=\alpha_{2cH,em}(\omega_{a},\omega_{b})+\alpha_{2cH,abs}(\omega_{a},\omega_{b}),$ (104)